OpenStudy (anonymous):

(Un) a sequence U0=1 ; Un+1=(Un-1)/2 1) alpha is real number =/0 we have Vn=Un+alpha What is the alpha value that make Vn geometric sequence

OpenStudy (anonymous):

\[U_{0}=1 ; U_{n+1}=\frac{ U_{n}-1}{ 2}\] \[V_{n}=U_{n}+alpha\]

terenzreignz (terenzreignz):

For it to be a geometric sequence, it must be the case that \[\large U_{n+1} = kU_n\] For any nonzero constant, k.

OpenStudy (anonymous):

what about our case

terenzreignz (terenzreignz):

processing... this is tricky, lol

OpenStudy (anonymous):

:) Please if you now something try to write it mathematically

terenzreignz (terenzreignz):

Do you have to add alpha to U[0] as well?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the value of alpha to make vn geometric

terenzreignz (terenzreignz):

So both sequences start at 1?

OpenStudy (anonymous):

vn linked by Un , i think

OpenStudy (anonymous):

@sirm3d what do you want to say

OpenStudy (sirm3d):

geometric sequence \(V_n\) \[V_{n+1}=rV_n\] substituting \[U_{n+1}+\alpha=r(U_n+\alpha\\\frac{U_n-1}{2}+\alpha=rU_n+r\alpha\]

OpenStudy (sirm3d):

\[\frac{1}{2}U_n-\frac{1}{2}+\alpha=rU_n+r\alpha\] therefore \[r=\frac{1}{2}\] and \[-\frac{1}{2}+\alpha = \frac{1}{2}\alpha\] solving alpha, \[\alpha=1\]

OpenStudy (anonymous):

this is @sirm3d you are the best

OpenStudy (sirm3d):

thank you

terenzreignz (terenzreignz):

hang on...

terenzreignz (terenzreignz):

Yeah, it works o.O

OpenStudy (anonymous):

Thank you guys

OpenStudy (sirm3d):

yw

OpenStudy (sirm3d):

put \[\frac{V_{n+1}}{V_n}=\text{constant} = r\]

OpenStudy (anonymous):

ok ok ok oko k

OpenStudy (anonymous):

\[\frac{ U_{n}-1+2\alpha }{ 2(U_{n}+\alpha)}=C\]

OpenStudy (anonymous):

If you let alpha by 1 you make un+1/2(un+1)=2 (constant)

OpenStudy (sirm3d):

clear the denominator first by writing it as \[U_n-1+2\alpha=2C(U_n+\alpha)\] expand the right hand side, then compare the coefficients of \(U_n\)

OpenStudy (anonymous):

C=1/2 ?

OpenStudy (sirm3d):

right. use that value to solve \(\alpha\)

OpenStudy (anonymous):

aaaaah yes

OpenStudy (anonymous):

but it is dificult to solve like this equations ha ?

OpenStudy (sirm3d):

not really. the key to solving the equation is to compare coefficients of \(U_n\)

OpenStudy (anonymous):

it is Conformity

OpenStudy (anonymous):

and now how to write Vn terms n

OpenStudy (anonymous):

The Vn term

OpenStudy (sirm3d):

write the terms of \(U_n\) and \(V_n\) \[U_0=1, V_0=1+1=2\\U_1=\frac{1-1}{2}=0, V_1=0+1=1\\U_2=\frac{0-1}{2}=-\frac{1}{2},V_2=-\frac{1}{2}+1=\frac{1}{2}\\V_n=2\left(\frac{1}{2}\right)^n\]